How many moles of HI are present in a 1000.0 g sample?
1
Expert's answer
2020-04-23T13:06:12-0400

Amount of substance (ν) can be found by the formula:


 ν = m/Mr,

where m is the mass of substance in grams,

Mr is the molecular mass of substance in grams/mole (can be found as a sum of the atomic mass of each elements - it can be found in the periodic table).


νHI = mHI/Mr(HI) = 1000 g / (127 + 1) g/mol = 1000 g/128 g/mol = 7,8125 mol ≈ 7,8 mol.


Answer: 7,8 mol.